# FedSRD: Sparsify-Reconstruct-Decompose for Communication-Efficient Federated Large Language Models Fine-Tuning

### v2

@@ 5.2.2. Abation Study
To validate each component of our FedSRD framework, we conducted an ablation study on Llama3.2-3B, with results in Table 5. Starting with the full model, we sequentially removed or replaced key components. Disabling the adaptive sparsity ratio (using a fixed ratio instead) causes a slight performance drop, affirming the benefit of dynamic adjustment. Replacing our importance-aware sparsification with a random baseline (Yu et al., 2024) leads to a significant degradation, highlighting that our decomposition method effectively preserves aggregated knowledge while reducing download communication costs. Further, ablating our alternative decomposition (reverting to dense SVD download) also harms performance, confirming its effectiveness in preserving knowledge while reducing download costs. Finally, removing the reconstruction step (reverting to naive LoRA matrix averaging) results in severe performance loss, proving that full-rank aggregation is crucial for mitigating client heterogeneity. These results confirm that each component plays an indispensable and synergistic role in achieving the dual objectives of high model performance and extreme communication efficiency. Additionally, we show in Appendix E that our importance-aware sparsification also works as a more effective plugin to enhance conventional methods like FedAvg.
@@ 5.3. Computation Time Analysis
We measured the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B models, with results reported in Table 7. Among the baselines, FedAvg provides a baseline for time consumption. FedProx requires slightly more time due to its proximal term, whereas FFA-LoRA is faster as it only trains and aggregates half of the LoRA matrices. FlexLoRA consumes more time performing SVD for matrix decomposition, and FedICU is the most time-intensive due to its computationally inefficient disentangled aggregation. For our methods, while the full FedSRD requires more time for its $r$-rank SVD projection, our efficient variant, FedSRD-e, introduces negligible additional computation compared to FedAvg, while exhibiting performance second only to the full FedSRD. This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLMs fine-tuning.
@@ Appendix A Pseudo-Code of FedSRD and FedSRD-e
2:Initial Synchronization: Each client $i\in\mathcal{C}$ sets its local state $B_{i}^{0}\leftarrow B^{0},A_{i}^{0}\leftarrow A^{0}$.
3:for each communication round $t=1,2,\dots,T$ do
4:  if $t>1$ then
5:   Server broadcasts $\Delta B_{s}^{t-1}$ or $\Delta A_{s}^{t-1}$.
6:   for each client $i\in\mathcal{C}$ in parallel do
7:     Client updates local model: $B_{i}^{t-1}\leftarrow B_{i}^{t-2}+\Delta B_{s}^{t-1}$, $A_{i}^{t-1}\leftarrow A_{i}^{t-2}+\Delta A_{s}^{t-1}$.
8:  for each client $i\in\mathcal{C}$ in parallel do
9:   $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}\leftarrow\text{ClientTraining}(i)$
10:  Server-Side Reconstruction and Aggregation:
11:  for each client $i\in\mathcal{C}$ do
12:   Client reconstructs its new state $B_{i}^{t},A_{i}^{t}$ using Eq. 4
13:   Server computes client’s full-rank matrix: $W_{i}^{t}\leftarrow B_{i}^{t}A_{i}^{t}$.
14:  Aggregate in full-rank space: $W^{t}\leftarrow\frac{1}{|\mathcal{C}|}\sum_{i}W_{i}^{t}$.
15:  Server-Side Alternating Decomposition:
16:  $W^{t}_{r}\leftarrow\text{SVD}_{r}(W^{t})$. $\triangleright$ FedSRD-e bypasses this SVD step.
17:  Let $\Delta W^{t}_{update}=W^{t}_{r}-B^{t-1}A^{t-1}$. $\triangleright$ For FedSRD-e, use $W^{t}$ instead of $W^{t}_{r}$.
18:  if $t$ is even then
19:   $\Delta A^{t}\leftarrow(B^{t-1})^{\dagger}(\Delta W^{t}_{update})$
20:   $\Delta B^{t}\leftarrow\mathbf{0}$
21:  else
22:   $\Delta B^{t}\leftarrow(\Delta W^{t}_{update})(A^{t-1})^{\dagger}$
23:   $\Delta A^{t}\leftarrow\mathbf{0}$
24:  Update global state on server: $B^{t}\leftarrow B^{t-1}+\Delta B^{t}$, $A^{t}\leftarrow A^{t-1}+\Delta A^{t}$.
25:  Sparsify for download: $\Delta B_{s}^{t},\Delta A_{s}^{t}\leftarrow\text{DARE}(\Delta B^{t},\Delta A^{t})$.
26:
27:function ClientTraining(ii)
28:  Fine-tune model on local data $\mathcal{D}_{i}$ to get new $B^{\prime}_{i},A^{\prime}_{i}$.
29:  Compute local updates: $\Delta B_{i}^{t}\leftarrow B^{\prime}_{i}-B_{i}^{t-1}$, $\Delta A_{i}^{t}\leftarrow A^{\prime}_{i}-A_{i}^{t-1}$.
30:  Importance-Aware Sparsification:
31:  Calculate importance scores $I(\Delta B_{i}^{t}),I(\Delta A_{i}^{t})$ using Eq. 6.
32:  Calculate adaptive sparsity ratios $\rho_{\Delta B_{i}^{t}},\rho_{\Delta A_{i}^{t}}$.
33:  Prune updates to get sparse $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ adaptively.
34:  return $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ to server.
@@ Appendix B Rationale of the Importance Score Calculation
Appendix C Description of Baselines
Our baselines span from foundational FL algorithms to advanced techniques specifically designed for communication-efficient and effective federated fine-tuning of large language models.
FedAvg (McMahan et al., 2017) is the first algorithm in federated learning. It establishes a simple yet effective paradigm for collaborative model training. In each communication round, clients download current global model parameters. Each client then performs multiple steps of local training on its private dataset. The resulting updated model weights are sent back to the server, which aggregates them by computing a weighted average to produce the global model for the subsequent round. FedAvg serves as the fundamental baseline for performance and communication cost.
FedProx (Li et al., 2020) is a key extension to FedAvg designed to address the challenge of statistical heterogeneity (non-IID data) across clients. Non-IID data can cause local client models to drift significantly from the global consensus, leading to convergence instability. FedProx mitigates this by introducing a proximal term to the local objective function on each client. This term regularizes the local training process by penalizing the squared Euclidean distance between the local model weights and the global model weights, thereby constraining local updates and promoting more stable convergence.
FFA-LoRA (Sun et al., 2024) is a parameter-efficient fine-tuning method tailored for the federated setting, focusing on reducing communication overhead and accurate aggregation. It utilizes the Low-Rank Adaptation (LoRA) strategy. To achieve communication efficiency, FFA-LoRA freezes the randomly initialized LoRA $A$ matrices across all clients and throughout the training process. Consequently, clients only need to train and communicate the $B$ matrices, drastically reducing the number of parameters transmitted in each round.
FlexLoRA (Bai et al., 2024) proposes a more sophisticated aggregation mechanism for federated LoRA fine-tuning. It posits that directly averaging the low-rank matrices from different clients may be suboptimal. Instead, clients upload their locally trained LoRA matrices. The server first reconstructs the full-rank weight updates for each client ($B_{i}A_{i}$) and averages them. Subsequently, the server performs Singular Value Decomposition (SVD) to find a new, optimal low-rank approximation, which is then decomposed into new $B$ and $A$ matrices for client downloading.
FedICU (Liao et al., [n. d.]) is an advanced FL method that decomposes the global model into consensus and divergence components to handle heterogeneity better. The core mechanism is a consensus-divergence splitting performed on the server-side during aggregation. Consensus components and divergence components of all clients are aggregated separately. FEdICU further proposes the selective updating strategy that uploads only more specialized parameters for aggregation.
Fed-DARE addresses the communication bottleneck in FL by integrating the DARE (Drop and Rescale) technique (Yu et al., 2024). In this approach, clients sparsify their local LoRA updates by randomly dropping parameters. The remaining non-zero elements are then rescaled to preserve the update’s expected magnitude, resulting in highly sparse matrices that significantly reduce the communication payload. We adopt a sparsity ratio of 0.9 as recommended by DARE. As shown in Figure 2, this random-based approach outperforms magnitude-based sparsification, while higher sparsity ratios lead to performance degradation. This justifies our choice and the exclusion of magnitude-based methods as a baseline.
@@ Appendix C Description of Baselines
Appendix D Communication Cost Analysis
@@
Appendix E Importance-aware Sparsification Suits Well for FedAvg
@@
Appendix F Justification for Equation 7
@@
Appendix G Magnitude of $B^{t-1},A^{t-1}$ and $\Delta B^{t},\Delta A^{t}$
@@
Appendix H Performance of Base Models and Locally Trained Models

### v3

@@ 5.2.2. Abation Study
To validate each component of our FedSRD framework, we conducted an ablation study on Llama3.2-3B, with results in Table 4. Starting with the full model, we sequentially removed or replaced key components. Disabling the adaptive sparsity ratio (using a fixed ratio instead) causes a slight performance drop, affirming the benefit of dynamic adjustment. Replacing our importance-aware sparsification with a random baseline (Yu et al., 2024) leads to a significant degradation, highlighting that our decomposition method effectively preserves aggregated knowledge while reducing download communication costs. Further, ablating our alternative decomposition (reverting to dense SVD download) also harms performance, confirming its effectiveness in preserving knowledge while reducing download costs. Finally, removing the reconstruction step (reverting to naive LoRA matrix averaging) results in severe performance loss, proving that full-rank aggregation is crucial for mitigating client heterogeneity. These results confirm that each component plays an indispensable and synergistic role in achieving the dual objectives of high model performance and extreme communication efficiency. Additionally, we show in Appendix D that our importance-aware sparsification also works as a more effective plugin to enhance conventional methods like FedAvg.
@@ 5.3. Computation Time Analysis
We report the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B in Table 6. Among the baselines, FedAvg provides a baseline for time consumption. FedProx requires slightly more time due to its proximal term, whereas FFA-LoRA is faster as it only trains and aggregates half of the LoRA matrices. FlexLoRA consumes more time performing SVD for matrix decomposition, and FedICU is the most time-intensive due to its computationally inefficient disentangled aggregation. For our methods, while the full FedSRD requires more time for its $r$-rank SVD projection, our efficient variant, FedSRD-e, introduces negligible additional computation compared to FedAvg, while exhibiting performance second only to the full FedSRD. This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLMs fine-tuning.
@@ 6. Conclusion
Acknowledgements.
This work was supported by the National Key R&D Program of China under Grant No. 2022YFB2703301. This work was also supported by the Science and Technology Development Program of Two Districts in Xinjiang, China under Grant No. 2024LQ03004.
@@ Appendix A Pseudo-Code of FedSRD and FedSRD-e
2:for each communication round $t=1,2,\dots,T$ do
3:  if $t>1$ then
4:   Server broadcasts $\Delta B_{s}^{t-1}$ or $\Delta A_{s}^{t-1}$.
5:   for each client $i\in\mathcal{C}$ in parallel do
6:    Client updates local model: $B_{i}^{t-1}\leftarrow B_{i}^{t-2}+\Delta B_{s}^{t-1}$, $A_{i}^{t-1}\leftarrow A_{i}^{t-2}+\Delta A_{s}^{t-1}$.
7:   end for
8:  end if
9:  for each client $i\in\mathcal{C}$ in parallel do
10:   $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}\leftarrow\text{ClientTraining}(i)$
11:  end for
12:  Server-Side Reconstruction and Aggregation:
13:  for each client $i\in\mathcal{C}$ do
14:   Client reconstructs its new state $B_{i}^{t},A_{i}^{t}$ using Eq. 4
15:   Server computes client’s full-rank matrix: $W_{i}^{t}\leftarrow B_{i}^{t}A_{i}^{t}$.
16:  end for
17:  Aggregate in full-rank space: $W^{t}\leftarrow\frac{1}{|\mathcal{C}|}\sum_{i}W_{i}^{t}$.
18:  Server-Side Alternating Decomposition:
19:  $W^{t}_{r}\leftarrow\text{SVD}_{r}(W^{t})$. {FedSRD-e bypasses this SVD step.}
20:  Let $\Delta W^{t}_{update}=W^{t}_{r}-B^{t-1}A^{t-1}$. {For FedSRD-e, use $W^{t}$ instead of $W^{t}_{r}$.}
21:  if $t$ is even then
22:   $\Delta A^{t}\leftarrow(B^{t-1})^{\dagger}(\Delta W^{t}_{update})$
23:   $\Delta B^{t}\leftarrow\mathbf{0}$
24:  else
25:   $\Delta B^{t}\leftarrow(\Delta W^{t}_{update})(A^{t-1})^{\dagger}$
26:   $\Delta A^{t}\leftarrow\mathbf{0}$
27:  end if
28:  Update global state: $B^{t}\leftarrow B^{t-1}+\Delta B^{t}$, $A^{t}\leftarrow A^{t-1}+\Delta A^{t}$.
29:  Sparsify for download: $\Delta B_{s}^{t},\Delta A_{s}^{t}\leftarrow\text{DARE}(\Delta B^{t},\Delta A^{t})$.
30:end for
31:
32:Function ClientTraining(ii):
33:  Fine-tune model on local data $\mathcal{D}_{i}$ to get new $B^{\prime}_{i},A^{\prime}_{i}$.
[…]
35:  Importance-Aware Sparsification:
36:  Calculate importance scores $I(\Delta B_{i}^{t}),I(\Delta A_{i}^{t})$ using Eq. 6.
37:  Calculate adaptive sparsity ratios $\rho_{\Delta B_{i}^{t}},\rho_{\Delta A_{i}^{t}}$.
38:  Prune updates to get sparse $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ adaptively.
39:return $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ to server.
@@ Appendix B Rationale of the Importance Score Calculation
B.1. Theoretical Analysis of Pruning Error
In this section, we provide a theoretical justification for the proposed importance score. We demonstrate that our greedy pruning strategy, based on the importance score defined in Equation 6, is mathematically equivalent to minimizing the upper bound of the reconstruction error of the full-rank weight update $\Delta W$.
B.1.1. Error Bound Derivation
We analyze the approximation error for the term $\Delta W_{B}=\Delta B_{i}^{t}A_{i}^{t-1}$ (the derivation for $\Delta A$ is symmetric). Let $\mathcal{S}$ denote the set of indices in $\Delta B_{i}^{t}$ selected for pruning. The pruning error matrix $\mathcal{E}$ is the sum of the rank-1 contributions of the removed elements:
(20) $\mathcal{E}=\sum_{(u,v)\in\mathcal{S}}(\Delta B_{i}^{t})_{uv}\cdot(\mathbf{e}_{u}\otimes\mathbf{a}_{v}^{T}),$
where $\mathbf{a}_{v}^{T}$ is the $v$-th row of $A_{i}^{t-1}$. By applying the Triangle Inequality to the Frobenius norm of $\mathcal{E}$, we derive the error upper bound:
(21) $\|\mathcal{E}\|_{F}$ $=\left\|\sum_{(u,v)\in\mathcal{S}}(\Delta B_{i}^{t})_{uv}(\mathbf{e}_{u}\otimes\mathbf{a}_{v}^{T})\right\|_{F}$
$\leq\sum_{(u,v)\in\mathcal{S}}\underbrace{|(\Delta B_{i}^{t})_{uv}|\cdot\|\mathbf{a}_{v}\|_{2}}_{I(\Delta B_{i}^{t}[u,v])}.$
The right-hand side is exactly the sum of the importance scores of the pruned parameters. Thus, pruning elements with the lowest scores $I(\cdot)$ strictly minimizes the upper bound of the weight update reconstruction error $\|\Delta W\|_{F}$.
B.1.2. Comparison with Gradient-based Pruning
Gradient-based pruning methods are impractical for Federated LoRA Fine-tuning for several reasons:
  • Computational and Memory Overhead: Calculating the exact influence of parameters on the loss usually requires second-order derivatives (e.g., Hessian). For LLMs, computing or even approximating the Hessian is prohibitively expensive for resource-constrained edge devices. Even first-order methods require retaining gradients, doubling the memory footprint during the scoring phase.
  • Federated Constraints: In a federated setting, data is non-IID. A gradient-based score computed on local data may not accurately reflect the parameter’s importance to the global model, potentially leading to overfitting local distributions.
In contrast, our proposed importance score is a data-free and zero-overhead metric. It relies solely on the magnitude of the weights and the structural correlations within the LoRA matrices ($\Delta B$ and $A$), which are already present in memory. This ensures that our method minimizes the deviation in the weight space ($\|\Delta W\|$) efficiently, offering a superior trade-off between communication efficiency and model performance.
@@ B.1.2. Comparison with Gradient-based Pruning
Appendix C Communication Cost Analysis
@@
Appendix D Importance-aware Sparsification Suits Well for FedAvg
@@
Appendix E Justification for Equation 7
@@
Appendix F Magnitude of $B^{t-1},A^{t-1}$ and $\Delta B^{t},\Delta A^{t}$
@@
Appendix G Performance of Base Models and Locally Trained Models